# Quantifying and Mitigating Unimodal Biases in Multimodal Large Language Models: A Causal Perspective

### v3

@@ Abstract
Recent advancements in Large Language Models (LLMs) have facilitated the development of Multimodal LLMs (MLLMs). Despite their impressive capabilities, MLLMs often suffer from an over-reliance on unimodal biases (e.g., language bias and vision bias), leading to incorrect answers in complex multimodal tasks. To investigate this issue, we propose a causal framework to interpret the biases in Visual Question Answering (VQA) problems. Within our framework, we devise a causal graph to elucidate the predictions of MLLMs on VQA problems, and assess the causal effect of biases through an in-depth causal analysis. Motivated by the causal graph, we introduce a novel MORE dataset, consisting of 12,000 VQA instances. This dataset is designed to challenge MLLMs’ abilities, necessitating multi-hop reasoning and the surmounting of unimodal biases. Furthermore, we propose two strategies to mitigate unimodal biases and enhance MLLMs’ reasoning capabilities, including a Decompose-Verify-Answer (DeVA) framework for limited-access MLLMs and the refinement of open-source MLLMs through fine-tuning. Extensive quantitative and qualitative experiments offer valuable insights for future research. Our project page is at https://opencausalab.github.io/MORE.
@@
Meiqi Chen [^1] This work was done during her internship at Shanghai AI Laboratory. Peking University meiqichen@stu.pku.edu.cn                        Yixin Cao Singapore Management University caoyixin2011@gmail.com
Yan Zhang [^2] Corresponding author. Peking University zhyzhy001@pku.edu.cn                        Chaochao Lu [^2] Corresponding author. Shanghai AI Laboratory luchaochao@pjlab.org.cn
@@ 1 Introduction
Following the success of Large Language Models (LLMs) Ouyang et al. (2022); Touvron et al. (2023b), Multimodal LLMs (MLLMs) OpenAI (2023); Team et al. (2023) have been proposed for various vision-language tasks Fu et al. (2023); Liu et al. (2023b). Despite their promising results, it remains unclear if they truly understand images and text in the context of multi-modal reasoning.
@@
In order to investigate the issue of MLLMs’ Over-REliance (MORE) on such unimodal biases, we propose a causal framework to interpret and quantify language and vision biases. Specifically, we first define the causal graph of MLLM’s prediction on VQA problems. The causal graph is constructed based on a variety of causal factors that are integral to the prediction process, such as image and question text. Then, we identify a set of interventions in the context of VQA problems, thereby ascertaining the causal effect of unimodal biases on the prediction capabilities of MLLMs via $do$-calculus Pearl (1995). By quantifying such causal effects, we can evaluate MLLM’s sensitivity and robustness against unimodal biases.
@@
Based on the above causal analysis, we curate a novel dataset termed MORE, comprising 12,000 VQA instances. This dataset advances existing VQA datasets by introducing a dedicated evaluation of unimodal biases. To facilitate the evaluation, we adopt a Multiple Choice Question (MCQ) format, where each instance consists of an image, a question, and four candidate options. The image is sourced from an existing VQA dataset Chen et al. (2023). For question and option curation, we incorporate a knowledge graph (KG) Wang et al. (2021), allowing us to better simulate MLLMs to navigate the corresponding spurious paths within the causal graph. Specifically, the options consist of one correct answer, and three distractors targeting language bias, vision bias, and multi-hop reasoning, respectively. We also provide the reasoning path, designated as causal rationale, in the KG for each instance, offering interpretability for evaluation. As summarized in Table 1, compared to existing VQA datasets, MORE features external knowledge, multi-hop reasoning, unimodal bias evaluation, and rationale, showing better comprehensiveness. Experimental results on six leading MLLMs reveal that: 1) most MLLMs perform much poorly on MORE, showing an obvious tendency to rely on unimodal biases. 2) MLLMs still struggle to achieve precise semantic understanding when dealing with multimodal reasoning.
Furthermore, to mitigate the unimodal biases and enhance MLLM’s reasoning abilities, we also propose two strategies. The basic idea is to leverage step-by-step multimodal reasoning to mitigate spurious paths. The first strategy involves prompt engineering, featuring limited-access MLLMs (e.g., GPT-4V OpenAI (2023)) through a Decompose-Verify-Answer (DeVA) framework to explicitly guide MLLMs to extract information from multiple modals comprehensively. The second strategy requires fine-tuning open-source MLLMs such as LLaVA Liu et al. (2023a) based on our dataset, instructing them with the generated causal rationale.
Overall, our main contributions are as follows:
  • We propose a causal framework with a well-defined causal graph to interpret and quantify the biases in VQA problems.
  • We construct a new MORE dataset, designed to challenge MLLMs to overcome reliance on unimodal biases via multi-hop causal reasoning, demonstrating superior comprehensiveness compared to existing VQA datasets.
  • We propose two strategies to mitigate the unimodal biases and enhance the reasoning ability of current MLLMs, including prompting and fine-tuning ways.
@@ 2 A Causal Framework
In this section, inspired by Stolfo et al. (2022), we first introduce the causal graph of MLLM’s prediction on VQA problems. Then, we leverage the causal graph to elucidate the biases inherent in VQA, especially vision and language biases. Finally, we assess the causal effects of such biases on MLLMs’ prediction by performing controlled interventions Pearl (1995).
@@ 2.1 Problem Setup
We consider an entity-centric VQA problem, denoted by $M$, consisting of a question $Q$ and an image $I$. The image depicts a specific entity and the question is related to the entity. The question $Q:=(S,T)$ is composed of two distinct elements: the core semantic content ${S}$, which conveys the genuine intent of the question, and the textual surface form ${T}$, which is irrelevant to the question’s core meaning. The image $I:=(E,C)$, instead, contains the core entity of interest $E$ and the irrelevant visual context $C$. The model’s final answer/prediction is represented by $A$. In this paper, we use lowercase letters to represent an instance of its corresponding capital variable.
2.2 Causal Graph of MLLMs’ Prediction
Inspired by the intuitive reasoning mechanisms observed in human cognition Zellers et al. (2019); Fei et al. (2022); Stolfo et al. (2022), we formulate the causal mechanisms underlying human problem-solving in a VQA problem $m$:
${s}=f_{c_{1}}({q}),\ {e}=f_{c_{2}}({i}),\ g=f_{c_{3}}({s,e}),$ (1)
where the cognitive process $f_{c_{1}}$ is employed to decipher the core semantic meaning $s$ of the question $q$, and $f_{c_{2}}$ is employed to extract the core entity $e$ from the image $i$. Then, the function $f_{c_{3}}$ is utilized to combine $s$ and $e$, yielding the final result $g$. These mechanisms are depicted within the green subgraph $G_{h}$ of Figure 2.
In contrast, the process of a model to tackle the same VQA problem $m$ involves the following causal mechanisms:
$a=f_{b}({q},{i}),\vspace{-5pt}$ (2)
where $a$ represents the model’s prediction. The function $f_{\text{b}}$ operates as a black box, leaving uncertainty regarding which aspects of $q$ and $i$ are considered by the model and how they interact with each other.
To further analyze how a model tackles the VQA problem, we draw all possible causal mechanisms that might take place in the complete causal graph in Figure 2. Notable causal mechanisms include:
@@ 2.2 Causal Graph of MLLMs’ Prediction
The model may directly process the question $Q$ in two ways: by focusing on the core semantics $S$ via the causal path ${Q}\rightarrow S\rightarrow A$, or on the irrelevant part $T$ via the causal path ${Q}\rightarrow T\rightarrow A$. Both pathways lead to language bias, e.g., the focus on “the next World Cup ” in Figure 1 (a).
@@
The essence of correct reasoning lies in the model’s grasp of the underlying causal mechanisms necessary for solving a VQA problem. As illustrated in Figure 2), the green subgraph $G_{h}$ embodies this comprehension. It implies an understanding of how both the image and the question jointly influence the ground-truth result ${G}$ (via ${E}\rightarrow G$ and ${S}\rightarrow G$). Consequently, the model’s predictions should exhibit sensitivity and robustness to changes in the ground truth Stolfo et al. (2022), namely ${G}\rightarrow A$. No spurious correlations should directly affect $A$ unless it goes through the mediator $G$. Based on the above analyses, we articulate the concept of model sensitivity and robustness on VQA problems:
  • Sensitivity, which assesses if the model adjusts its predictions properly when the correct answer changes, i.e., $A$ responds to changes in $G$.
  • Robustness, which assesses the undesired direct causal effect of unimodal biases, e.g., $C\rightarrow A$ and $T\rightarrow A$, where lower effects indicate greater robustness to input changes that do not alter the correct answer.
@@ 2.3 Causal Analysis of VQA Biases
Once the desired causal mechanism is formulated and the path of unimodal biases is defined, we are able to quantify the causal effects of each causal factor on another by performing controlled interventions Pearl (1995).
Causal Interventions
In the context of VQA, we adopt the following interventions to quantify the causal effects of images and questions on the model’s predictions:
  • 1).
    Interventions on $Q$. The question $Q$ can be modified in two ways: (i) altering both $S$ and $T$ , or (ii) altering $T$ but keeping $S$ unaffected.
  • 2).
    Interventions on $I$. The image $I$ can be replaced with an alternative image $I^{\prime}$ in three ways: (i) altering both $E$ and $C$ , or (ii) altering $C$ but keeping $E$ unaffected, or (iii) altering $E$ but keeping $C$ unaffected.
Note that we do not solely alter $S$ within $Q$, because it is not feasible to intervene on the core semantics $S$ of a question without affecting the surface text $T$ of it.
@@
Next, we explain how we can obtain the causal effect from the interventions. Consider an intervention formulated as $\operatorname{do}\left(X:x\rightarrow x^{\prime}\right)$, where $X\in\{I,E,C,Q,T\}$ is a factor in the context of the VQA problem defined by $M=\{I,Q\}$. For simplicity, the probability distribution prior to the intervention, $\mathbb{P}(A\mid{I},{Q})$, is denoted by $P$, while the distribution following the intervention is represented as $P^{\prime}$. Following Pearl (1995), we measure the causal effect of each factor in our causal graph by employing the distance metric $\delta$ to gauge the divergence between distributions $P$ and $P^{\prime}$:
$\operatorname{CE}=\delta\left(P,P^{\prime}\right).$ (3)
It is important to note that $\operatorname{CE}$ denotes the causal effect, which can refer to either of the two forms: the total causal effect ($\operatorname{TCE}$), signifying the joint effect across all causal paths from one variable to another; and the direct causal effect ($\operatorname{DCE}$), indicating the effect of the directed causal path devoid of intermediary variables Pearl (2022).
Following  Stolfo et al. (2022), we quantify the causal effect of factor $X$ on the model’s answer $A$ by assessing the change in the predicted result:
$\delta_{\mathrm{cp}}\left(P,P^{\prime}\right):=\mathbb{I}\left(a\neq a^{\prime}\right)$ (4)
where $a=\arg\max_{x}P(x),\text{ and }a^{\prime}=\arg\max_{x}P^{\prime}(x)$, $\mathbb{I}$ denotes an indicator of the “change answer” event.
Causal Effects of the Questions
For the total causal effect of the question on the answer, we can quantify it by intervening on $Q$:
@@
Note that, this $\operatorname{TCE}$ contains two different types of paths that show how $Q$ affects $A$, as illustrated in Figure 2:
  • The desired decision-making route we aim for the model to adopt ( the ${Q}\rightarrow S\rightarrow G\rightarrow A$ path), where the model responds accordingly to alterations in the ground truth answer.
  • The spurious correlation that the model might have learned (e.g., the ${Q}\rightarrow T\rightarrow A$ path, where the model depends on certain linguistic contexts potentially linked to their prevalence in the training corpus).
Instead of changing the core semantic meaning, maintaining $S$ constant will enable us to derive the $\operatorname{DCE}$ of the textual surface form $T$ on $A$:
@@
It should be noted that the $\operatorname{DCE}({T}\rightarrow A$) equates to the $\operatorname{TCE}$ of $T$ on $A$, due to there is no mediator between $T$ and $A$. Besides, as discussed in “Causal Interventions”, solely intervening on $S$ without influencing the textual surface $T$ is unattainable. However, comprehending the causal impact of $S$ on $A$ becomes feasible through the comparison of two known quantities: $\operatorname{TCE}({Q}\text{ on }A)$ and $\operatorname{DCE}({T}\rightarrow A)$.
Causal Effects of the Images
It is evident from Figure 2 that the causal structure of images is exactly the same as that of questions. Therefore, the total causal effect of the image $I$ on the answer can be directly obtained in a similar manner, i.e.,
@@
Likewise, maintaining $E$ constant during each intervention on $I$ allows us to quantify the $\operatorname{DCE}$ of the irrelevant visual context $C$ on $A$, which refers to the strength of the path ${C}\rightarrow A$:
@@
Maintaining $C$ and $G$ constant during each intervention on $I$ instead allows us to quantify $\operatorname{DCE}({E}\rightarrow A)$:
@@
Note that in practical situations, it is significantly challenging to intervene on the core entity $E$ while simutaniously controlling for the irrelevant visual context $C$, the ground truth answer $G$, and the question $Q$. Therefore, this type of intervention will not be considered in subsequent experiments.
Overall, calculating $\operatorname{TCE}$ helps us assess a model’s sensitivity (response to changes in ground truth), while $\operatorname{DCE}$ evaluates its robustness (stability of predictions against spurious correlations with fixed ground truth).
3 Constructing the MORE Dataset
@@
Based on the causal analysis in Section 2, we construct a novel MORE dataset that necessitates MLLMs to transcend unimodal biases and thoroughly integrate information from both text and images to select the correct answer. To ensure quality, we employ various evaluations to analyze the quality of the generated data. The data generation process is illustrated in Figure 3.
3.1 Images and Knowledge Graph Collection
We begin with an existing VQA dataset, INFOSEEK Chen et al. (2023), which links entities depicted in images to information sourced from Wikipedia, requiring a VQA model to answer a question about the associated entity. For instance in Figure 3, the entity associated with the showcased vehicle is “Chrysler PT Cruiser”, with the question being “What equipment or engine is used by this vehicle to provide power?” In this example, terms such as “Chrysler PT Cruiser”, “this vehicle”, and the relevant region in the image all refer to the same entity.
Based on the images and corresponding entity information, we identify all $n$-order neighbors ($n\in\{1,2\}$) related to the associated entity in a knowledge graph (KG) - Wikidata5M Wang et al. (2021), which is built upon Wikipedia data.
3.2 Subgraph Sampling
Motivated by the causal analysis discussed in Section 2, we aim to construct multi-hop queries that require overcoming unimodal biases to be answered correctly. To this end, we first identify a subgraph of an entity and its $n$-order neighbors in the KG. We then filter paths in this subgraph that meet two criteria: 1) Uniqueness of Paths: the path from the associated entity to the selected neighbor is unique, and 2) Shared-Type Relations: they share a same-type relation pointing to a unique entity, and these two pointed entities are not the same. The first criterion guarantees the uniqueness of the correct answer, while the second introduces interference and challenges MLLM’s reasoning ability.
For instance, in the subgraph of Figure 3, “Fiat 500X” is a first-order neighbor of “Chrysler PT Cruiser”, uniquely linked by the “followed by” relation. These two entities are linked to different entities “The Chrysler Corporation” and “Fiat Automobiles S.p.A” through the same type of relations “brand”. Meeting these criteria, we form the multi-hop query: “Chrysler PT Cruiser $\stackrel{{\scriptstyle\text{followed by}}}{{\longrightarrow}}$iat 500X $\stackrel{{\scriptstyle\text{brand}}}{{\longrightarrow}}$ Fiat Automobiles S.p.A”, with “Chrysler PT Cruiser” as the start and “Fiat Automobiles S.p.A” as the answer. In this way, other entities of the filtered subgraph may serve as distractors to challenge MLLMs’ reasoning ability, which will be discussed in the next section.

### v4

@@ Abstract
Recent advancements in Large Language Models (LLMs) have facilitated the development of Multimodal LLMs (MLLMs). Despite their impressive capabilities, MLLMs often suffer from over-reliance on unimodal biases (e.g., language bias and vision bias), leading to incorrect answers or hallucinations in complex multimodal tasks. To investigate this issue, we propose a causal framework to interpret the biases in Visual Question Answering (VQA) problems. Within this framework, we conduct an in-depth causal analysis to assess the causal effect of these biases on MLLM predictions. Based on the analysis, we introduce 1) a novel MORE dataset with 12,000 challenging VQA instances requiring multi-hop reasoning and overcoming unimodal biases. 2) a causality-enhanced agent framework CAVE that guides models to comprehensively integrate information from different modalities and mitigate biases. Our experiments show that MLLMs perform poorly on MORE, indicating strong unimodal biases and limited semantic understanding. However, when integrated with our CAVE, promising improvements in reasoning and bias mitigation can be seen. These findings provide important insights for the development of more robust MLLMs and contribute to the broader goal of advancing multimodal AI systems capable of deeper understanding and reasoning. Our project page is at https://github.com/OpenCausaLab/MORE.
@@
Meiqi Chen1,2111This work was done during her internship at Shanghai AI Laboratory., Yixin Cao3, Yan Zhang1,$2$Corresponding author., Chaochao Lu4222Corresponding author. 1State Key Laboratory of General Artificial Intelligence, Peking University, Beijing, China 2School of Intelligence Science and Technology, Peking University 3School of Computer Science, Fudan University 4Shanghai Artificial Intelligence Laboratory meiqichen@stu.pku.edu.cn, caoyixin2011@gmail.com, zhyzhy001@pku.edu.cn, luchaochao@pjlab.org.cn
@@ 1 Introduction
Following the success of Large Language Models (LLMs) Ouyang et al. (2022); Touvron et al. (2023b), Multimodal LLMs (MLLMs) OpenAI (2023); Team et al. (2023) have been proposed for various vision-language tasks Fu et al. (2023); Liu et al. (2023b). Despite their promising results, it remains unclear if they truly understand images and text in the context of multimodal reasoning.
@@
To investigate the issue of MLLMs’ Over-REliance (MORE) on such unimodal biases, we propose a causal framework to interpret and quantify language and vision biases. We begin by defining a causal graph of MLLM’s prediction on VQA problems, built on key causal factors like images and questions. Then, we identify a set of interventions in the context of VQA problems, thereby ascertaining the causal effect of unimodal biases on MLLM predictions via $do$-calculus Pearl (1995). This allows us to evaluate the sensitivity and robustness of MLLMs against unimodal biases.
@@
Based on the above causal analysis, we curate a novel dataset termed MORE, comprising 12,000 VQA instances. This dataset advances existing VQA datasets by introducing a dedicated evaluation of unimodal biases. We adopt a Multiple Choice Question (MCQ) format to facilitate the evaluation, where each instance consists of an image, a question, and four candidate options. The image is sourced from an existing VQA dataset Chen et al. (2023). For question and option curation, we incorporate a knowledge graph (KG) Wang et al. (2021), allowing us to simulate MLLMs to navigate potential paths within the causal graph. Specifically, the options consist of one correct answer, and three distractors targeting language bias, vision bias, and multi-hop reasoning, respectively. We also provide the reasoning path, designated as causal rationale, in the KG for each instance, offering interpretability for evaluation. As summarized in Table 1, compared to existing VQA datasets, MORE features better comprehensiveness.
Furthermore, motivated by the causal analysis, we propose CAVE, a causality-enhanced method to mitigate unimodal biases in MLLMs. CAVE encompasses a diverse set of workflows, including question decomposition, causality-based enhanced self-reflection, external knowledge retrieval, and answer verification. This framework guides models in explicitly and comprehensively integrating information from multiple modalities while helping to prevent biased decision-making and the selection of incorrect shortcuts.
Through extensive experiments on MORE with several leading MLLMs, we observe that: 1) most MLLMs perform much poorly on MORE, showing strong reliance on unimodal biases and low robustness to disturbances; 2) MLLMs still struggle to achieve precise semantic understanding in multimodal reasoning; 3) With our propose method CAVE, we can mitigate the unimodal bias in MLLMs, yet it still falls short of the ideal. This indicates that addressing the unimodal bias in MLLMs is a highly challenging issue that merits further exploration. Overall, our main contributions are as follows:
  • We propose a causal framework to interpret and quantify the biases in VQA problems.
  • We construct a new dataset, MORE, which requires multi-hop reasoning and overcoming biases, demonstrating superior comprehensiveness to existing VQA datasets.
  • We conduct extensive experiments on MORE and propose a causality-enhanced method CAVE to mitigate the unimodal biases, providing insights for future work.
@@ 2 A Causal Framework
Inspired by Stolfo et al. (2023), we introduce a causal graph for MLLM predictions on VQA problems, highlighting language and vision biases and assessing their causal effects via controlled interventions Pearl (1995).
@@ 2.1 Problem Setup
We consider an entity-centric VQA problem, $M$, involving a question $Q:=(S,T)$ and an image $I:=(E,C)$. Here, $S$ is the core semantic content of $Q$, and $T$ is the textual form unrelated to $Q$’s main meaning. $I$ includes the primary entity $E$ and surrounding visual context $C$. The model outputs an answer, $A$. Notation: lowercase letters denote instances of corresponding uppercase variables.
2.2 Causal Graph of MLLM Predictions
Inspired by human cognitive processes Zellers et al. (2019); Stolfo et al. (2023), we formulate the causal mechanisms in human problem-solving for a VQA problem $m$: ${s}=f_{c_{1}}({q}),\ {e}=f_{c_{2}}({i}),\ g=f_{c_{3}}({s,e})$. This involves decoding the question $q$ to extract its semantic meaning $s$ through cognitive process $f_{c_{1}}$, and identifying the core entity $e$ from the image $i$ through $f_{c_{2}}$. These are combined by $f_{c_{3}}$ to produce the result $g$, as shown in the green subgraph $G_{h}$ of Figure 2. In contrast, the model’s approach to the same VQA problem employs $a=f_{b}({q},{i})$, a black-box function where it is unclear how question and image inputs are utilized and interact to form the prediction $a$. For deeper analysis, we illustrate possible causal mechanisms in Figure 2. Notable mechanisms include:
@@ 2.2 Causal Graph of MLLM Predictions
The model may directly process the question $Q$ in two ways: by focusing on the core semantics $S$ via ${Q}\rightarrow S\rightarrow A$, or on the irrelevant part $T$ via ${Q}\rightarrow T\rightarrow A$. Both pathways lead to language bias, e.g., the focus on “the next World Cup ” in Figure 1 (a).
@@
Correct reasoning in VQA problems hinges on understanding causal mechanisms, as depicted in Figure 2. The subgraph $G_{h}$ represents comprehension of how image and question jointly affect the ground-truth result ${G}$, through ${E}\rightarrow G$ and ${S}\rightarrow G$. This understanding should lead to model predictions that are sensitive and robust to changes in ${G}$, namely ${G}\rightarrow A$, with no spurious effects on $A$ unless it passes the mediator $G$. Model performance could therefore be evaluated by its: 1) Sensitivity, which assesses how well the model adapts to changes in the correct answer, i.e., $A$ responds to changes in $G$. 2) Robustness, which measures resistance to unimodal biases, e.g., $C\rightarrow A$ and $T\rightarrow A$, where minimal bias effects imply higher robustness to input changes that do not alter the ground-truth answer.
@@ 2.3 Causal Analysis of VQA Biases
We adopt the controlled interventions as outlined by Pearl (1995) to quantify the causal effects of questions and images on model predictions.
Causal Interventions for VQA
1) Interventions on $Q$. The question $Q$ can be modified in two ways: (i) altering both $S$ and $T$ , or (ii) altering $T$ but keeping $S$ unaffected. 2) Interventions on $I$. The image $I$ can be replaced with an alternative image $I^{\prime}$ in three ways: (i) altering both $E$ and $C$ , or (ii) altering $C$ but keeping $E$ unaffected, or (iii) altering $E$ but keeping $C$ unaffected. Note that we do not solely alter $S$ within $Q$, because it is not feasible to intervene on the core semantics $S$ of a question without affecting the surface text $T$ of it.
@@
We assess the causal effects using $\operatorname{do}\left(X:x\rightarrow x^{\prime}\right)$, where $X\in\{Q,T,I,C,E\}$ is a factor in the VQA problem $M$. The pre-intervention probability distribution $\mathbb{P}(A\mid I,Q)$ is denoted as $P$, and the post-intervention distribution as $P^{\prime}$. Following Pearl (1995), the causal effect is evaluated using a distance metric $\delta$, as $\operatorname{CE}=\delta(P,P^{\prime})$, where $\operatorname{CE}$ denotes the causal effect. It can refer to 1) the total causal effect ($\operatorname{TCE}$), signifying the joint effect across all causal paths from one variable to another; or 2) the direct causal effect ($\operatorname{DCE}$), indicating the effect of the directed causal path devoid of intermediary variables Pearl (2022). Following Stolfo et al. (2023), we assess the causal effect of factor $X$ on the model’s answer $A$ by comparing the change in predicted results, $\delta_{\mathrm{cp}}(P,P^{\prime}):=\mathbb{I}(a\neq a^{\prime})$, where $a=\arg\max_{x}P(x)$ and $a^{\prime}=\arg\max_{x}P^{\prime}(x)$, with $\mathbb{I}$ indicating a change in the answer.
Causal Effects of Questions
We assess $\operatorname{TCE}$ of a question $Q$ on an answer $A$ by intervening on $Q$:
@@
This $\operatorname{TCE}$ contains two different types of paths that show how $Q$ affects $A$, as illustrated in Figure 2: 1) The intended decision-making pathway: ${Q}\rightarrow S\rightarrow G\rightarrow A$, responding to changes in the ground truth. 2) Potential spurious correlations, e.g., ${Q}\rightarrow T\rightarrow A$, where the model may depend on certain linguistic patterns from training data.
Maintaining the core semantics $S$ constant, we can assess $\operatorname{DCE}$ of the textual surface $T$ on $A$:
@@
As discussed in “Causal Interventions for VQA”, solely intervening on $S$ without altering $T$ is impractical. However, Understanding $S$’s causal impact on $A$ is feasible by comparing two known quantities: $\operatorname{TCE}({Q}\text{ on }A)$ and $\operatorname{DCE}({T}\rightarrow A)$.
Causal Effects of Images
The causal structure of images mirrors that of questions, as shown in Figure 2. Thus, we can derive $\operatorname{TCE}$ of image $I$ on the answer as follows:
@@
Likewise, maintaining $E$ constant during each intervention on $I$ allows us to quantify the $\operatorname{DCE}$ of the irrelevant visual context $C$ on $A$:
@@
Maintaining $C$ and $G$ constant during each intervention on $I$ instead allows us to quantify the $\operatorname{DCE}$ of the core entity $E$ on $A$:
@@
Overall, calculating $\operatorname{TCE}$ helps us assess models’ sensitivity (response to changes in ground truth), while $\operatorname{DCE}$ evaluates its robustness (stability of predictions against spurious correlations).
3 the MORE Dataset
@@
From Section 3.1 to 3.3, we construct the MORE dataset that exploits the unimodal biases defined in Section 2.2 and requires multi-hop reasoning. In Section 3.4, we discuss how to quantify the causal effects of images and questions (defined in Section 2.3) using our constructed dataset.
3.1 Preparatory Steps
Images and Knowledge Graph Collection
We begin with an existing VQA dataset, INFOSEEK Chen et al. (2023), which links image entities to Wikipedia information, requiring a VQA model to answer related questions. For instance in Figure 3, the entity “Chrysler PT Cruiser” prompts the question, “What equipment or engine is used by this vehicle to provide power?” Here, terms like “Chrysler PT Cruiser” and “this vehicle” all refer to the same entity in the image. We then identify all $n$-order neighbors ($n\in\{1,2\}$) of the associated entity within a knowledge graph (KG), Wikidata5M Wang et al. (2021), which is built upon Wikipedia data.
Subgraph Sampling
[…]